OpenStudy (mathisfun13):

convert from radian measure to degree measure pi/6

OpenStudy (zzr0ck3r):

radians*(180/pi)=degrees

OpenStudy (mathisfun13):

1080 but it said it was wrong

OpenStudy (zzr0ck3r):

(pi/6)*(180/pi) = ?

OpenStudy (mathisfun13):

i divide 3.14/6?

OpenStudy (jhannybean):

\[\large\frac{\pi}{6}\cdot\frac{180}{\pi}=?\]

OpenStudy (mathisfun13):

30

OpenStudy (jhannybean):

goodjob.

OpenStudy (mathisfun13):

thanks(: